# A Generalized Control Function Approach to Production Function Estimation

### v2

@@ 2 Results
A complication arises because estimating $\tilde{\varphi}_{it}$ and $\tilde{q}_{it}-\tilde{f}_{it}$ requires itself a plugin estimator for $\text{E}[q_{it-1}|x_{it-1}]$. To the best of our knowledge, the literature has not yet developed a treatment for such “double plugin” estimators. In what follows, we therefore focus on the case where the control function simplifies from $h\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$ to $h\left(z_{it}\setminus z^{\dagger}_{it}\right)$ as in Example 1. Thus, the plugin estimator for $\text{E}[q_{it-1}|x_{it-1}]$ is no longer required, and $\tilde{\varphi}_{it}=\text{E}\left[\left.\varphi(z_{it})\right|z_{it}\setminus z^{\dagger}_{it}\right]$ and $\tilde{q}_{it}-\tilde{f}_{it}=\text{E}\left[\left.q_{it}-f(k_{it},v_{it})\right|z_{it}\setminus z^{\dagger}_{it}\right]$.
While our generalized control function approach accommodates the joint evolution of productivity $\omega_{it}$ and latent demand shocks $\delta_{it}$, it is not costless. If moment condition (1) holds, then for any choice of the special instrument $z^{\dagger}_{it}\subsetneq z_{it}\setminus(k_{it-1},v_{it-1})$ it is less efficient than the OP/LP/ACF procedure. This reflects a general tradeoff between robustness and efficiency in nonparametric estimation. Our generalized control function approach makes this tradeoff explicit by allowing a researcher to target greater robustness (smaller $z^{\dagger}_{it}$) or greater efficiency (larger $z^{\dagger}_{it}$).
@@ 3 Monte Carlo Exercise
We use GMM estimation based on moment condition (5) to estimate the production function parameters $\theta=\left(\alpha,\rho,\nu\right)$. Our instruments are $z_{it}=\left(k_{it},k_{it-1},v_{it-1},p_{it-1},p^{V}_{it},p^{V}_{it-1}\right)$ and our special instrument is $z^{\dagger}_{it}=p^{V}_{it}$ as in Example 1. Our weighting function $\varphi(z_{it})$ is the complete set of Hermite polynomials of total degree 4 in the variables in $z_{it}$. We estimate $\tilde{\varphi}_{it}=\text{E}\left[\varphi(z_{it})|z_{it}\setminus z^{\dagger}_{it}\right]$ by OLS using the complete set of Hermite polynomials of total degree $d$ in the variables in $z_{it}\setminus z^{\dagger}_{it}$. We proceed similarly to estimate $\tilde{q}_{it}-\tilde{f}_{it}=\text{E}\left[q_{it}-f(k_{it},v_{it})|z_{it}\setminus z^{\dagger}_{it}\right]$. The latter must be re-estimated at each iteration of the GMM problem. Even though the control function $h\left(z_{it}\setminus z^{\dagger}_{it}\right)=\text{E}\left[\omega_{it}+\varepsilon_{it}|z_{it}\setminus z^{\dagger}_{it}\right]$ is absent from moment condition (5), the total degree $d$ implicitly determines how well we can approximate it. We accordingly explore $d\in\{2,3,4,5\}$. We provide further details on the GMM estimator in the Supplemental Appendix.
@@ Results.
Turning to the generalized control function approach in this paper, Figure 1 shows the distribution of the average log markup. As can be seen, the results rapidly improve with the total degree $d$ of the complete set of Hermite polynomials in the variables in $z_{it}\setminus z^{\dagger}_{it}$. The bias decreases from $0.0507$ for $d=2$ to 0.0006-0.0006 for $d=4$ and the mean squared error from to $0.0032$ to $0.0007$. There are no further improvements going from $d=4$ to $d=5$.
@@ 4 Concluding Remarks
We provide conditions for consistently estimating the production function in empirically relevant settings that are outside the scope of the OP/LP/ACF procedure. As such, our approach complements the OP/LP/ACF framework. Our approach generalizes the control function that is already present in OP and requires solving a straightforward GMM problem. We hope that it proves valuable for applied researchers seeking to estimate the production function and the markup from it.
@@
Johns Hopkins University November 27, 2025
@@
Let $f^{0}(k_{it},v_{it})$ denote the true production function. Because $q_{it}-f^{0}(k_{it},v_{it})=\omega_{it}+\varepsilon_{it}$, Assumption 1 implies that $f^{0}(k_{it},v_{it})$ satisfies moment condition (3) for the control function $h^{0}\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)=\text{E}\left[\left.q_{it}-f^{0}(k_{it},v_{it})\right|z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right]$. Let $\tilde{f}(k_{it},v_{it})$ be some production function that also satisfies moment condition (3) for some control function $\tilde{h}\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$. To show that $\frac{\partial f(k_{it},v_{it})}{\partial v_{it}}$ is nonparametrically point-identified, we show that, for any $k_{it}$, the difference $f^{0}(k_{it},v_{it})-\tilde{f}(k_{it},v_{it})$ does not change with $v_{it}$ almost surely.
@@
$=$ $\text{E}\left[\left.\tilde{h}\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)-h^{0}\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)\right|z_{it},\text{E}[q_{it-1}|x_{it-1}]\right]$
$=$ $\tilde{h}\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)-h^{0}\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right).$
@@
$=$ $\text{E}\left[\left.f^{0}(k_{it},v_{it})-\tilde{f}(k_{it},v_{it})\right|z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right].$
@@
$\text{E}\left[\left.\kappa\left(v_{it},z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)\right|z_{it},\text{E}[q_{it-1}|x_{it-1}]\right]$
$=$ $\text{E}\left[\left.\kappa\left(v_{it},z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)\right|z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right]$
holds only if $\kappa$ does not change with $v_{it}$ almost surely. Recalling that $k_{it}\in z_{it}\setminus z^{\dagger}_{it}$ and setting $\kappa\left(v_{it},z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)=f^{0}(k_{it},v_{it})-\tilde{f}(k_{it},v_{it})$ therefore establishes the result.
@@
Because $\varphi(z_{it})-\tilde{\varphi}_{it}$ is pinned down by $\left(z_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$, we have
@@
Next, we prove that moment condition (5) implies moment condition (4) with $h(z_{it}\setminus z^{\dagger}_{it}$, $\text{E}[q_{it-1}|x_{it-1}])=\tilde{q}_{it}-\tilde{f}_{it}$. Note that
@@
$=$ $\text{E}\left[\tilde{\varphi}_{it}\text{E}\left[\left.q_{it}-f(k_{it},v_{it})-\left(\tilde{q}_{it}-\tilde{f}_{it}\right)\right|z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right]\right]$
@@
where the first equality is due to the law of iterated expectations and the second equality to $\tilde{q}_{it}-\tilde{f}_{it}=\text{E}[q_{it}-f(k_{it},v_{it})|z_{it}\setminus{z^{\dagger}_{it}},\text{E}[q_{it-1}|x_{it-1}]]=0$ by the definition of $\tilde{q}_{it}-\tilde{f}_{it}$. Adding moment condition (5) and moment condition (4) implies moment condition (4) with $h\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)=\tilde{q}_{it}-\tilde{f}_{it}$.
Finally, we prove that moment condition (5) is Neyman orthogonal. Define the shorthand $\tilde{z}_{it}=\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$. Define $\mathcal{F}$ to be the set of function tuples $(\eta,\zeta)$ such that $\text{E}[\eta^{2}(\tilde{z}_{it})]<\infty$ and $\text{E}[\zeta^{2}(\tilde{z}_{it})]<\infty$. Because $\varphi(z_{it})$ and $q_{it}-f(k_{it},v_{it})$ have finite $L^{2}$ norms, for any $(\eta,\zeta)\in\mathcal{F}$ and for any $\lambda=(\lambda_{1},\lambda_{2})\in^{2}$, we have
$\text{E}\left|\left(\varphi(z_{it})-\tilde{\varphi}_{it}-\lambda_{1}\eta(\tilde{z}_{it})\right)\left(q_{it}-f(k_{it},v_{it})-\left(\tilde{q}_{it}-\tilde{f}_{it}+\lambda_{2}\zeta(\tilde{z}_{it})\right)\right)\right|<\infty.$
@@
$\left.\frac{\partial\text{E}(\varphi(z_{it})-\tilde{\varphi}_{it}-\lambda_{1}\eta(\tilde{z}_{it}))(q_{it}-f(k_{it},v_{it})-(\tilde{q}_{it}-\tilde{f}_{it}+\lambda_{2}\zeta(\tilde{z}_{it})))}{\partial\lambda}\right|_{\lambda=0}$
$=$ $\begin{pmatrix}\text{E}\left[\eta(\tilde{z}_{it})\left(q_{it}-f(k_{it},v_{it})-\left(\tilde{q}_{it}-\tilde{f}_{it}\right)\right)\right]\\ -\text{E}\left[(\varphi(z_{it})-\tilde{\varphi}_{it})\zeta(\tilde{z}_{it})\right]\end{pmatrix}=\begin{pmatrix}0\\ 0\end{pmatrix}$,
@@
Recall that $\tilde{\varphi}_{it}=\text{E}\left[\varphi(z_{it})|z_{it}\setminus z^{\dagger}_{it}\right]$. To the extent that terms in $\varphi(z_{it})$ can be perfectly predicted by the complete set of Hermite polynomials of total degree $d$ in the variables in $z_{it}\setminus z^{\dagger}_{it}$, the matrix $\Phi$ with rows $\varphi(z_{it})-\tilde{\varphi}_{it}$ is rank deficient. In solving the GMM problem, we therefore use a selection of columns from the matrix $\Phi$ that has full rank. To construct this selection, we start with an empty matrix and keep adding columns from the matrix $\Phi$ as long as this increases the rank.

### v3

@@ 2 Results
A complication arises because estimating $\tilde{\varphi}_{it}$ and $\tilde{q}_{it}-\tilde{f}_{it}$ requires itself a plugin estimator for $\text{E}[q_{it-1}|x_{it-1}]$. To the best of our knowledge, the literature has not yet developed a treatment for such “double plugin” estimators. In what follows, we therefore focus on the case where the control function simplifies from $h\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$ to $h\left(z^{c}_{it}\right)$ as in Example 2. Thus, the plugin estimator for $\text{E}[q_{it-1}|x_{it-1}]$ is no longer required, and $\tilde{\varphi}_{it}=\text{E}\left[\left.\varphi(z_{it})\right|z^{c}_{it}\right]$ and $\tilde{q}_{it}-\tilde{f}_{it}=\text{E}\left[\left.q_{it}-f(k_{it},v_{it})\right|z^{c}_{it}\right]$.
While our generalized control function approach accommodates the joint evolution of productivity $\omega_{it}$ and latent demand shocks $\delta_{it}$, it is not costless. If moment condition (1) holds, then for any choice of the special instrument $z^{s}_{it}\subsetneq z_{it}\setminus(k_{it-1},v_{it-1})$ it is less efficient than the OP/LP/ACF procedure. This reflects a general tradeoff between robustness and efficiency in nonparametric estimation. Our generalized control function approach makes this tradeoff explicit by allowing a researcher to target greater robustness (smaller $z^{s}_{it}$) or greater efficiency (larger $z^{s}_{it}$).
@@ 3 Monte Carlo Exercise
We use GMM estimation based on moment condition (5) to estimate the production function parameters $\theta=\left(\alpha,\rho,\nu\right)$. Our instruments are $z_{it}=\left(k_{it},k_{it-1},v_{it-1},p_{it-1},p^{V}_{it},p^{V}_{it-1}\right)$ and our special instrument is $z^{s}_{it}=p^{V}_{it}$ as in Example 2. Our weighting function $\varphi(z_{it})$ is the complete set of Hermite polynomials of total degree 4 in the variables in $z_{it}$. We estimate $\tilde{\varphi}_{it}=\text{E}\left[\varphi(z_{it})|z^{c}_{it}\right]$ by OLS using the complete set of Hermite polynomials of total degree $d$ in the variables in $z^{c}_{it}$. We proceed similarly to estimate $\tilde{q}_{it}-\tilde{f}_{it}=\text{E}\left[q_{it}-f(k_{it},v_{it})|z^{c}_{it}\right]$. The latter must be re-estimated at each iteration of the GMM problem. Even though the control function $h\left(z^{c}_{it}\right)=\text{E}\left[\omega_{it}+\varepsilon_{it}|z^{c}_{it}\right]$ is absent from moment condition (5), the total degree $d$ implicitly determines how well we can approximate it. We accordingly explore $d\in\{2,3,4,5\}$. We provide further details on the GMM estimator in the Supplemental Appendix.
@@ Results.
Turning to the generalized control function approach in this paper, Figure 1 shows the distribution of the average log markup. As can be seen, the results rapidly improve with the total degree $d$ of the complete set of Hermite polynomials in the variables in $z^{c}_{it}$. The bias decreases from $0.0507$ for $d=2$ to 0.0006-0.0006 for $d=4$ and the mean squared error from to $0.0032$ to $0.0007$. There are no further improvements going from $d=4$ to $d=5$.
@@ 4 Concluding Remarks
We provide conditions for consistently estimating the production function in empirically relevant settings that are outside the scope of the OP/LP/ACF framework. As such, our approach complements the OP/LP/ACF framework. It generalizes the control function that is already present in the OP/LP/ACF framework and requires solving a straightforward GMM problem. We hope that it proves valuable for applied researchers seeking to estimate the production function and the markup from it.
@@
Johns Hopkins University December 9, 2025
@@
Let $f^{0}(k_{it},v_{it})$ denote the true production function. Because $q_{it}-f^{0}(k_{it},v_{it})=\omega_{it}+\varepsilon_{it}$, Assumption 1 implies that $f^{0}(k_{it},v_{it})$ satisfies moment condition (3) for the control function $h^{0}\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)=\text{E}\left[\left.q_{it}-f^{0}(k_{it},v_{it})\right|z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right]$. Let $\tilde{f}(k_{it},v_{it})$ be some production function that also satisfies moment condition (3) for some control function $\tilde{h}\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$. To show that $\frac{\partial f(k_{it},v_{it})}{\partial v_{it}}$ is nonparametrically point-identified, we show that, for any $k_{it}$, the difference $f^{0}(k_{it},v_{it})-\tilde{f}(k_{it},v_{it})$ does not change with $v_{it}$ almost surely.
@@
$=$ $\text{E}\left[\left.\tilde{h}\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)-h^{0}\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)\right|z_{it},\text{E}[q_{it-1}|x_{it-1}]\right]$
$=$ $\tilde{h}\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)-h^{0}\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right).$
@@
$=$ $\text{E}\left[\left.f^{0}(k_{it},v_{it})-\tilde{f}(k_{it},v_{it})\right|z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right].$
@@
$\text{E}\left[\left.\kappa\left(v_{it},z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)\right|z_{it},\text{E}[q_{it-1}|x_{it-1}]\right]$
$=$ $\text{E}\left[\left.\kappa\left(v_{it},z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)\right|z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right]$
holds only if $\kappa$ does not change with $v_{it}$ almost surely. Recalling that $k_{it}\in z^{c}_{it}$ and setting $\kappa\left(v_{it},z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)=f^{0}(k_{it},v_{it})-\tilde{f}(k_{it},v_{it})$ therefore establishes the result.
@@
This is because moment condition (3) implies $h(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}])=\text{E}[q_{it}-f(k_{it},v_{it})|z_{it},\text{E}[q_{it-1}|x_{it-1}]]$ so that the control function has to satisfy $h(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}])=\text{E}[q_{it}-f(k_{it},v_{it})|z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]]=\tilde{q}_{it}-\tilde{f}_{it}$. Because $\varphi(z_{it})-\tilde{\varphi}_{it}$ is pinned down by $\left(z_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$, we have
@@
Next, we prove that moment condition (5) implies moment condition (4) with $h\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)=\tilde{q}_{it}-\tilde{f}_{it}$. Note that
@@
$=$ $\text{E}\left[\tilde{\varphi}_{it}\text{E}\left[\left.q_{it}-f(k_{it},v_{it})-\left(\tilde{q}_{it}-\tilde{f}_{it}\right)\right|z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right]\right]$
@@
where the first equality is due to the law of iterated expectations and the second equality to $\tilde{q}_{it}-\tilde{f}_{it}=\text{E}[q_{it}-f(k_{it},v_{it})|z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]]=0$ by the definition of $\tilde{q}_{it}-\tilde{f}_{it}$. Adding moment condition (5) and moment condition (4) implies moment condition (4) with $h\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)=\tilde{q}_{it}-\tilde{f}_{it}$.
Finally, we prove that moment condition (5) is Neyman orthogonal. Define the shorthand $\tilde{z}^{c}_{it}=\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$. Define $\mathcal{F}$ to be the set of function tuples $(\eta,\zeta)$ such that $\text{E}[\eta^{2}(\tilde{z}^{c}_{it})]<\infty$ and $\text{E}[\zeta^{2}(\tilde{z}^{c}_{it})]<\infty$. Because $\varphi(z_{it})$ and $q_{it}-f(k_{it},v_{it})$ have finite $L^{2}$ norms, for any $(\eta,\zeta)\in\mathcal{F}$ and for any $\lambda=(\lambda_{1},\lambda_{2})\in^{2}$, we have
$\text{E}\left|\left(\varphi(z_{it})-\tilde{\varphi}_{it}-\lambda_{1}\eta(\tilde{z}^{c}_{it})\right)\left(q_{it}-f(k_{it},v_{it})-\left(\tilde{q}_{it}-\tilde{f}_{it}+\lambda_{2}\zeta(\tilde{z}^{c}_{it})\right)\right)\right|<\infty.$
@@
$\left.\frac{\partial\text{E}(\varphi(z_{it})-\tilde{\varphi}_{it}-\lambda_{1}\eta(\tilde{z}^{c}_{it}))(q_{it}-f(k_{it},v_{it})-(\tilde{q}_{it}-\tilde{f}_{it}+\lambda_{2}\zeta(\tilde{z}^{c}_{it})))}{\partial\lambda}\right|_{\lambda=0}$
$=$ $\begin{pmatrix}\text{E}\left[\eta(\tilde{z}^{c}_{it})\left(q_{it}-f(k_{it},v_{it})-\left(\tilde{q}_{it}-\tilde{f}_{it}\right)\right)\right]\\ -\text{E}\left[(\varphi(z_{it})-\tilde{\varphi}_{it})\zeta(\tilde{z}^{c}_{it})\right]\end{pmatrix}=\begin{pmatrix}0\\ 0\end{pmatrix}$,
@@
Recall that $\tilde{\varphi}_{it}=\text{E}\left[\varphi(z_{it})|z^{c}_{it}\right]$. To the extent that terms in $\varphi(z_{it})$ can be perfectly predicted by the complete set of Hermite polynomials of total degree $d$ in the variables in $z^{c}_{it}$, the matrix $\Phi$ with rows $\varphi(z_{it})-\tilde{\varphi}_{it}$ is rank deficient. In solving the GMM problem, we therefore use a selection of columns from the matrix $\Phi$ that has full rank. To construct this selection, we start with an empty matrix and keep adding columns from the matrix $\Phi$ as long as this increases the rank.